OpenStudy (anonymous):

Use Euler's Formula to find the missing number. Vertices: 13 Edges: 26 Faces: ? 14 15 16 17

OpenStudy (anonymous):

Euler Formula for faces, vertices and edges F+V−E=2

OpenStudy (anonymous):

I think it would be 15

OpenStudy (anonymous):

i think your right

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Euler's formula states: r = m - v + 2, where r = faces, v = vertices, and m = edges r = 26 - 13 + 2 r = 15

OpenStudy (anonymous):

Thank you annas!

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Find the surface area of the cylinder to the nearest whole number. The figure is not drawn to scale. 905 in^2 1,282 in^2 2,187^2 7,691^2

OpenStudy (anonymous):

@annas can do this one i dont remember these haha

OpenStudy (anonymous):

surface area of cylinder is: 2 π r^2 + 2 π r h (2 * 3.1415 * (12)^2) + (2 * 3.1415 * 12 * 17)

OpenStudy (anonymous):

My answer wasn't even on there haha

OpenStudy (anonymous):

A = 1281.7 in^2 = 1282 in^2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Welcome ^.^

OpenStudy (anonymous):

Mind helping me with one or two more?

OpenStudy (anonymous):

I can try :)

OpenStudy (anonymous):

Find the surface area of the regular pyramid shown to the nearest whole number. The figure is not drawn to scale. 351m^2 415m^2 542m^2 790m^2

OpenStudy (anonymous):

hmm not very sure about it's formula :/ sowwy

OpenStudy (anonymous):

It's okay :) Thank you for helping me with the other one :)

OpenStudy (anonymous):

Welcome again :) & sowwy again for not able to help on this one :/